# UltraRay: Introducing Full-Path Ray Tracing in Physics-Based Ultrasound Simulation

### v2

@@ 1 Introduction
  • Code Publicly Available: We make the simulator’s code and scenes publicly available to facilitate future research and development.
@@ 2.3.2 Ray Traversal
$\omega_{r}=\omega_{i}+2\cos\theta_{r})\mathbf{n}$ (9)
$\omega_{t}=\eta\omega_{r}+(\eta\cos(\theta_{r}-\cos\theta_{t})$ (10)
@@
$A_{r}=\frac{Z_{1}\cos\theta_{r}-Z_{2}\cos\theta_{t}}{Z_{1}\cos\theta_{t}+Z_{2}\cos\theta_{r}}$ (11)
and due to energy conservation, the amplitude of the transmitted wave is $A_{t}=1.0-A_{r}$. In the Monte Carlo sampling scheme, we have to decide whether the ray reflects or transmits. In order to maintain a physically consistent reflected-to-transmitted ratio, we sample a variable $y$ from a uniform distribution $y\sim\mathcal{U}(0,1)$ and reflect if $y<A_{r}$.

### v3

@@ 1 Introduction
  • Code Availability: The simulator’s code and example scenes are openly accessible to support future research and development at
    https://github.com/Felixduelmer/UltraRay.
@@ 2.3.2 Ray Traversal
$\omega_{r}=\omega_{i}+2\cos\theta_{r}\mathbf{n}$ (9)
$\omega_{t}=\eta\omega_{i}+(\eta\cos\theta_{r}-\cos\theta_{t})\mathbf{n}$ (10)
@@
$A_{r}=\frac{Z_{1}\cos\theta_{r}-Z_{2}\cos\theta_{t}}{Z_{1}\cos\theta_{r}+Z_{2}\cos\theta_{t}}$ (11)
and the intensity of the reflected intensity is $I_{r}=I_{i}*(A_{r})^{2}$. Due to energy conservation, the following equation must hold: $I_{i}=I_{t}+I_{r}$. In the Monte Carlo sampling scheme, we have to decide whether the ray reflects or transmits. In order to maintain a physically consistent reflected-to-transmitted ratio, we sample a variable $y$ from a uniform distribution $y\sim\mathcal{U}(0,1)$ and reflect if $y<(A_{r})^{2}$.